OpenStudy (anonymous):

If h(x) = −5x − 10, find h(−2).

OpenStudy (anonymous):

it's not that hard.....y dont u try?

OpenStudy (unklerhaukus):

\[h(-2)=-5\times(-2)-10=\cdots\]

OpenStudy (anonymous):

-20 thks

OpenStudy (unklerhaukus):

no not -20

OpenStudy (unklerhaukus):

\[-5\times-2+-10=\cdots\neq-20\]

OpenStudy (unklerhaukus):

remember that a negative times a negative is a positive